Q9. The demand for widgets (Q) is given by the following equation Q 500 - 100 P. -50P, - 150 A, + 200A the price of widgets, currently at 25 P, the price of woozles, currently at 20 A,-Advertising on woozles currently at 10. A,-Advertising for widgets currently at 30 The cost per widget is currently 20 and the manufacturer where P behaves as a monopolist (a) What are current profits equal to? (b) Calculate the elasticity of demand for widgets with respect to its own price and the price of woozles. Provide a one sentence explanation for each elasticity measure (c) Calculate the elasticity of demand for widgets with respect to advertising on widgets and advertising on woozles. Provide a one-sentence explanation for each elasticity measure. (d) What are profits equal to if the manufacturer maximizes revenue? (e) What are profits equal to if the manufacturer maximizes profits? (f) Calculate the price elasticity of demand for widgets at the profit-maximizing price. (g) What is consumer surplus equal to at the profit-maximizing price?
Q8. The demand for Tesla cars (Q) is given by the following equation: Q 5480- 20 P 14P, +200P where P the price of Tesla cars, currently at 150. P, the price of Chrysler cars, currently at 60 P, the price of gasoline, currently at 1.4. The cost per Tesla car is currently 130 and the manufacturer behaves as a monopolist. (a) What are Tesla's current profits equal to? (b) Calculate the elasticity of demand for Tesla cars with respect to its own price, the price of Cals etei nes or onler and Tes cars substitutes or complements? What about Tesla cars and gasoline? (c) What are profits equal to if Tesla Motors maximizes revenue? (d) What are profits equal to if Tesla Motors maximizes profits? (e) Calculate the price elasticity of demand for Tesla cars at the profit-maximizing price. (f) What is consumer surplus equal to at the profit-maximizing price?
Answer #1

13b Qx (1So 1303660 Our Telsa 31ヌ8x 26 3660 3666 2 Scanned with o 83 amscanner06圣 1y 6 .233 за 20b 9, eems chryslt amd Talsa cans Scanned with CamScannerC. d T dP 2 d TR 0 -(16-130233 CamScanner20x 130 20 dPx = 230 40 23 o quantit7 -오-우nia 3 x Scanned with amscannerd Pa dpメ 230 2230 3 330 3o 66 50 s Scanned with CamScannery o 156 25-2a156 b) Bio elautii 160 -100 (155) Eoi, f , = 1-66 to SScanned wiEpx, By - CamScanner20メ ?y Try &メ 06 50 -150 Ay 266 Ax TR = 400メー0% dox 163 dTR Scanned wit CS CamScanne1163 R428 PX๑メ 2 &メー d頁 2 Qメ 26 ← 160 2-0 50 Cs Scanned with CamScanner-1m(.law = afァ.Ex 0ワ 30 5000 Scanned with CamScanner Cs